OpenStudy (anonymous):

find how many ways the integers 1,2,3,4,5,6,7,8 can be placed in a circle if: - Atleast three odd numbers are together?

OpenStudy (zarkon):

I get 2304

OpenStudy (zarkon):

\[(4!)^2+4\times3\times2\times4\times3\times3!=(4!)^2\times 4=2304\]

OpenStudy (anonymous):

why do you times it by 4 for?

OpenStudy (zarkon):

\[(4!)^2+4\times3\times2\times4\times3\times3!\] \[=(4!)^2+(4\times3\times2)\times4\times3!\times3\] \[=(4!)^2+4!\times4!\times3=(4!)^2+(4!)^2\times3=(4!)^2[1+3]\] \[=(4!)^2\times 4=2304\]